OpenStudy (anonymous):

how to show the work for 3(d+12)=8-4d

OpenStudy (anonymous):

Distribute like last time and then get d all by itself.

OpenStudy (anonymous):

help me plzzzz

OpenStudy (anonymous):

3(d + 12) = 8 - 4d 3d + 36 = 8 - 4d 3d = 8 - 36 - 4d 3d + 4d = 8- 36 7d = -28 d = -28/7 d = -4

OpenStudy (anonymous):

7(r+7)=5r+59